# Are LLMs Better than Reported? Detecting Label Errors and Mitigating Their Effect on Model Performance

### v1

@@ Acknowledgements
Ethics Statement
@@
Reproducibility Statement
In the study, we performed experiments involving human annotators, including both experts and crowd-workers. While the exact results of these experiments cannot be fully reproduced due to the inherent variability of human participants, the process can be replicated. To facilitate this, we provide detailed guidelines, platform setup, and technical specifications related to the annotation procedure of crowd-source (subsection 3.2, subsection A.1) and experts (subsection 3.2, subsection A.2).
For the experiments involving LLMs, we have included detailed procedural steps (subsection 2.2; subsection 3.2; and Appendix A.3), the prompts used (Figure 10), model versions (subsection 3.2; subsection 6.1; and Appendix C.2), hardware specifications and hyperparameter configurations (Appendix C.1), statistical measures (Appendix D), and evaluation metrics. These materials ensure that the LLM experiments are reproducible.
@@
Appendix A Annotation Approaches
A.1 Crowd-source
A.1.1 Annotation Procedure.
@@ A.1 Crowd-source
Each example was annotated by three annotators, that in addition to the binary label were requested to provide their confidence in their answer, and also write a short explanation for why they chose this label. Pre-qualifications included 50+ approved HITs, 97%+ approval rate, and locations from [USA, UK, Australia], which are all English-speaker countries. We disabled the possibility of right-click and Ctrl+c in the platform (as suggested by (Veselovsky et al., 2023a)), to prevent (as much as possible) the case where generative-AI (e.g., ChatGPT) will be applied to solve the task instead of humans solving it themselves (as shown by (Veselovsky et al., 2023b)). Maximum time allowed per HIT was 6 minutes, while the actual average execution time was 2:20 minutes for all assignments, and 3 minutes for approved assignments. The guidelines provided to annotators and the annotation platform layout are presented in Figure 8.
@@
Each annotation was manually reviewed and was rejected if the answers were not in line with the instructions, or if it was obvious that the task wasn’t done honestly. Overall, this task suffered from a high rejection rate of 49.2% (1163 rejected, 1200 approved). The main rejection reasons were: lack of meaningful explanation, obvious copy-paste annotations across different examples, explanation contradict with the label annotation, and cases where the explanation was a copy-paste of either the grounding or the statement.
@@
A.1.2 Consistency
Table 5: Distribution of crowd-source annotators. Each example was annotated by 3 workers. Plain segments are unanimous annotation, while dotted segments indicate examples where some annotators labeled as inconsistent, and other as consistent. For example, 19.8% of the examples had two inconsistent annotation, and one consistent annotation.
Table 6: How experts annotations have changed after the reconciliation phase. Most changes occur from 1 (consistent) to 0 (inconsistent).
Crowd workers showed near-random agreement, indicating relatively poor-quality annotations. Table 6 describes the distribution of annotations by MTurk workers. Only 40.8% from the examples were labeled unanimously, where the rest included annotations from both classes. In addition, if aggregating as majority vote, we get that 75.8% of the examples are labeled as consistent, which is far from the original distribution of classes. As mentioned before, even experts may miss a small inconsistency nuance, and finding it requires attention. Even from the subset of examples unanimously labeled as consistent, 37.9% have a label of inconsistent in both original and gold labels, which points at a lack of attention and thoroughness.
A.2 Experts
Experts annotation was using the platform of Label Studio. [^5] https://labelstud.io/ Layout design is presented in Figure 9. Examples were presented in random order, and neither the LLM prediction nor the original label were presented during the annotation. At the first stage, each example was annotated independently both experts. Afterwards, the human experts began in a second phase of a reconciliation, where a discussion was made over examples they disagree over. This reconciliation phase ended up with a much higher agreement, and higher-quality labels.
In the reconciliation phase, we observed that most changes (69.3%) were from label 1 to label 0, indicating that contradictions might be hard to find, and not all annotators catch them at first. For the full distribution of label change in the reconciliation phase, see Table 6.
@@ A.2 Experts
A.3 LLMs
We used four LLMs for the task of annotating a total of $160\times 4=640$ examples from four different datasets. Each model was run with four different prompts (see full prompts in Figure 10). We used a variety of terminology, as this task appear with different framing in different studies. For example, the premise-hypothesis terminology from classic NLI (MacCartney & Manning, 2009), or document-statement used in (Tam et al., 2023).
@@ A.3 LLMs
Appendix B Cost and Scalability
In section 5 we compare the different annotation approaches on label quality, consistency and cost-effectiveness. Here are the full details regarding run-time and costs.
In MTurk platform, a total of $400\times 3=1200$ annotations cost $572\$$, including 2 small pilot experiments. All annotations were prepared within a few hours. However, it demanded an additional and significant time for review, after which rejected examples returned to the pool. This annotation-review cycle was conducted for $\sim 5$ iterations. Inference via OpenAI’s API on GPT-4 cost $\sim 4.5\$$ per prompt. Inference via VertexAI’s API on PaLM2 cost $\sim 0.15\$$ per prompt. Both took $\sim 8$ minutes per prompt. Inference on Mistral and Llama3 was via the HuggingFace API, and its cost is estimated by the cost of using a suitable Virtual Machine (VM) on Google Cloud Platform (GCP) for the time of inference (1 minute per model)- $\sim 0.1\$$ per prompt. In total, annotating with an LLM is $\sim 10^{2}$ to $10^{3}$ times cheaper than a human annotator in a crowd-sourcing platform, and of course than human experts. It is also faster if considering reviewing procedure (excluding code writing or layout designing).
Appendix C Mislabeled Data Implications
C.1 Fine-tuning

### v2

@@ Acknowledgements
Limitations
While our study provides valuable insights into the role of LLMs in identifying label errors and improving dataset quality, several limitations should be considered. First, crowd workers encompass a broad range of annotators with varying expertise and training. Our analysis, focuses on the “common” crowd worker, typically an annotator selected with minimal qualifications, such as an approved task completion rate, and without specialized training. However, some datasets implement more selective strategies, such as requiring prior experience or task-specific training, which may yield more reliable labels. These "trained" crowd workers can be seen as an intermediate category between common annotators and experts, both in terms of cost and label quality. We chose to focus on the two endpoints, comparing common crowd workers and experts, to highlight clear contrasts in annotation quality and associated trade-offs. Importantly, we did not take crowd-worker annotations at face value; we applied filtering (based on the explanation crowd workers were asked to write for each example) to remove a substantial number of low-quality assignments, such as clearly invalid responses, in addition to enforcing minimal qualification criteria.
Second, our analysis does not account for potential data contamination, where LLMs may have been trained on the datasets we evaluate. However, since our analysis focuses on identifying and correcting label errors within these datasets, contamination would likely hinder rather than enhance our findings. If an LLM had memorized these datasets, it would be more likely to reproduce existing errors rather than detect and correct them, making contamination a potential limitation only for certain aspects of evaluation but not for our core claims.
Third, LLM-based annotations can vary depending on the choice of prompting strategies and ensemble methods. In this work, we use zero-shot prompting and simple averaging for ensembling. Still, alternative approaches – such as few-shot prompting, chain-of-thought reasoning (Wei et al., 2022), or self-refine (Madaan et al., 2023) – could improve annotation accuracy and consistency. Likewise, for ensembling, more advanced methods- such as percentile-based aggregation (Sherratt et al., 2023), error-aware weighting (Freund and Schapire, 1997), confidence-aware methods (Lee, 2010; Lu et al., 2024), or even LLM-based aggregation strategies like debate variants (Liang et al., 2023; Du et al., 2024) – may yield more reliable consensus labels. We leave the exploration of these strategies for future work and hope our study encourages such further research.
Finally, while our study does not cover the full range of NLP tasks, it is grounded in diverse and realistic labeling settings. The TRUE benchmark includes factual consistency annotations for summarization, dialogue, paraphrasing, and fact verification. SummEval adds ordinal labels and evaluates multiple dimensions of summary quality, such as fluency and coherence. These datasets differ in task framing, label format, and domain, providing a solid basis for analyzing label errors and their effects. Extending this analysis to other task types is a valuable direction for future work.
Ethical Considerations
@@
[…]
Appendix A Additional Experiments - SummEval
In addition to the datasets from the TRUE benchmark, we replicate our experiments on another dataset with a different objective and a different labeling scheme, to strengthen our results and conclusions.
A.1 Data
SummEval (Fabbri et al., 2021) is an extensive and commonly used summarization benchmark, evaluating the quality of multiple model-generated summarization outputs compared to a source CNN/DailyMail sources on four dimensions: coherence, relevance, consistency, and fluency. Each summarization is labeled on each dimension with five crowd-workers and three experts, enabling us to replicate some of the experiments without additional crowd-worker or expert annotation costs. The labeling schema is ordinal on a scale of 1 to 5 (higher is better). Note that this dataset does not have a singular gold-standard label per summarization, but rather a collection of annotations from experts and crowd-workers. Therefore, we will not claim to find label errors in this benchmark, but rather showcase our methodology as if the crowd-sourced annotations are the original labels for the dataset, and we have access to experts’ annotations for gold-standard reference, to determine if the LLM was correct when flagging examples.
A.2 Definitions
To apply our methods for error detection via LLMs ensemble, we first define the following:
Labels  We aggregate crowd-sourced annotations by their median, to construct a single original label on a scale of 1 to 5. Similarly, we take the median of the experts’ annotations to be a single gold-standard label.
A disagreement  We say that the LLM annotation disagrees with the original label if there is a difference of more than 1 between the scores. Smaller differences (e.g., 4 vs. 5) may reflect natural variation in subjective interpretation rather than a labeling mistake, and are therefore not considered strong disagreements. In practice, using a threshold of 1 results in over 50% of the dataset being flagged, making it difficult to isolate meaningful errors. We adopt this more conservative threshold to better reflect genuine annotation issues and reduce noise in our error detection process.
A.3 Experimental Setting
Similar to the description in subsection 4.2, we utilize two LLMs– GPT-4o (gpt-4o-2024-11-20) and Gemini 1.5 Flash (gemini-1.5-flash-002). We constructed four prompts, differing by phrasing and compatible with the four prompt template structures used for the TRUE benchmark experiments. The answer to each query was a JSON format with ’Relevance’, ’Coherence’, ’Consistency’, and ’Fluency’ as its keys. The scores are integers on a scale of 1 to 5, as are the ratings in the SummEval dataset. We extract the probability of each score possible through the log-probs for each score token. Finally, we average all models’ probabilities, to obtain an ensemble of LLMs, with $p$ being the distribution over the five possible scores.
A.4 Experiments and Results
A.4.1 Can LLMs Detect Label Errors?
We replicate the experiment described in subsection 5.2 with the appropriate adjustment for the SummEval dataset, based on the definitions above. The result is shown in Figure 2 (bottom). The plot presents the subset of examples where there was a disagreement between the crowd-sourced annotation and the LLMs’ annotation. Each bin represents the confidence of the LLMs in their predicted label. As there are five ordinal categories, even if there was a disagreement between two annotations, they both might be "wrong", where the expert’s answer is a third option. Therefore, to show clearer results, we do not resolve by experts "who is correct", but rather "who is more correct?". For completeness, we also provide the "both equally correct" option, for the case the expert’s label is exactly in the middle, and none is "more correct" than the other. The bins are relatively balanced in terms of the amount of examples per bin. Note that in contrast to the TRUE binary labeling scheme, where confidence 0.5 is the minimal threshold for an answer, here we start from 0.2.
From the results, we see a clear dominance of the LLM over the crowd-sourced annotations, for all confidence bins. This suggests that the LLMs not only detect error by flagging possibly mislabeled data points, but also provide better answers, which can account for error correction. Similar to the result on the TRUE benchmark, we observe a trend where as the LLMs’ confidence increases, they are more correct, indicating that they find label errors with higher precision. However, in this dataset, the difference from the original labels (in this case, the MTurk labels) is even more apparent, and the LLMs are correct even when with lower confidence.
A.4.2 The Power of Ensemble
We analyze the importance of utilizing more than a single model and a single prompt on two dimensions - performance compared to the gold labels (the quality of the annotations we utilize), and error detection (the ability to identify errors more accurately). For performance evaluation on the ordinal labels, we report Pearson correlation; for error detection evaluation, we report the F1-score based on binary error/not-error classification. See results in Figure 3 and discussion in LABEL:sec:_appendix-ensemble.
A.4.3 Annotation Approaches Comparison
In Appendix B, we thoroughly discuss the comparison between the different annotation approaches. For SummEval, experts and crowd-sourced annotations are provided. Together with our LLM-ensemble annotations (as described in subsection A.3), we analyze and compare the annotation approaches in terms of quality (see Figure 7 (bottom)) and consistency (see Table 5). To account for ordinal labels, we measure IAA via Krippendorff’s $\alpha$ (Krippendorff, 1970).
Appendix B Comparing Annotation Approaches
Our paper discusses three annotation approaches, each with its own benefits and drawbacks. These approaches differ in how they manage the trade-offs between label quality, scalability, and cost. In the following section, we discuss and compare their characteristics. A summary of this comparison is given in Figure 4.
B.1 Annotation Quality
Figure 7 (top) presents the F1-score between each pair of annotation approaches. As the figure shows, LLMs have disagreements with the original labels (0.72). Yet, as discussed in subsection 5.1, the original labels themselves contain mistakes, so this disagreement does not necessarily indicate poor performance of the LLMs. When considering the Gold as the true label, LLM performance increases to 0.83. This suggests that LLMs, despite their discrepancies with the original labels, perform closer to the truth than initially reported. The Gold label, obtained by experts, has high agreement with both the Original and LLM labels. On the other hand, the MTurk-Majority approach performs poorly, with near-random F1-scores compared to both the original and gold labels, and even when compared to its stricter variant, MTurk-Strict. The results indicate that basic crowd-sourcing, without additional training to enhance crowd-workers into specialized sub-experts, performs significantly worse compared to other approaches, including LLM-based methods. On the SummEval dataset (bottom of Figure 7), we observe similar results, where the LLMs are more correlated with the Experts rather than the crowd-workers, which in turn have almost-no-correlation with LLMs or experts’ annotations- this implies poor quality of the annotations obtained from crowd-sourcing. Still, we do not suggest that crowd-sourcing is inherently flawed; with proper task design and worker training, it may be suitable for certain subjective or human-centered tasks. However, we advocate for more careful consideration when using generic crowd annotations for evaluation.
Figure 7: Comparison between all annotation methods: (Top) on the TRUE benchmark, measured by the weighted-F1-score. Rows represent the "true" label and columns represent the "prediction". For instance, the score of LLMs compared to the Original label is 0.72. (Bottom) Comparison on the SummEval benchmark, measured by Pearson correlation (results are averaged over all dimensions).
Crowd-sourcing
For crowd-sourcing, the reported F1-score does not provide the complete picture. When we focus on individual annotators, we see that those who annotate more examples generally deliver higher-quality annotations, achieving greater accuracy when compared to both the original and gold labels (see Figure 5). This phenomenon can be explained by two hypotheses: (1) a learning process– as the annotators see more examples, they improve at the task, or (2) users who dedicate time to annotating multiple examples are likely those who either read the guidelines carefully and strive to perform the task to the best of their ability, or are naturally proficient at the task and therefore continue annotating. Even though annotators who label more instances tend to provide higher-quality annotations, they are less common—most annotators tend to stop after only a few examples. This distribution of annotators results in overall insufficient annotation quality. Pre-qualification tests are often used to shift this distribution from the "average worker" towards more experienced or dedicated annotators; however, this requires a significantly larger budget and greater micro-management involvement from the researcher.
B.2 Consistency
Usually, when annotating a dataset, more than one annotator is involved. This applies to crowd-workers, experts, and even LLMs- in this study, we use an ensemble of different LLMs and prompts. The use of multiple annotators, similar to an ensemble, is meant to overcome the variance between individuals, which can arise from the subjective nature of NLP tasks, different interpretations of instructions, lack of experience, task difficulty, and cognitive bias (Uma et al., 2021).
As such, a common practice in the NLP community is to report Inter Annotator Agreement (IAA)- a set of statistical measures used to evaluate the agreement between individuals. Typically, IAA can be viewed as an adjustment of the proportion of pairwise agreements, where 0.0 indicates random agreement. We focus on Fleiss’s $\kappa$ (Fleiss, 1971), as it accounts for label imbalance and multiple (>2>2) annotators. High IAA, or low variance between independent annotators, is considered an indicator of high-quality annotation. In Table Table 3, we report the agreement between annotators across different approaches. For LLMs, we report two variants: (1) same model, different prompts; and (2) different models, where each model’s result is the aggregation across prompts. For reference, we also include the IAA from the original annotations, as reported in the original papers: MNBM reported an average Fleiss’s $\kappa$ of 0.696 for the hallucination annotation task; BEGIN reported Krippendorff’s $\alpha$ (a generalization of Fleiss’s $\kappa$) of 0.7; VitaminC reported Fleiss’s $\kappa$ of 0.7065 on a sample of 2,000 examples; and PAWS reported a 94.7% agreement between a single annotator’s label and the majority vote on the Wikipedia subset used in TRUE.
Annotator group Krippendorff’s $\alpha$ %agreement #annotators Experts 0.584 60.4 3 MTurk [^2] These MTurk annotators were chosen with stricter pre-qualification criteria than those in the TRUE dataset and do not correspond to the MTurk line in the TRUE table. 0.496 65.6 5 LLM (different prompts) 4    GPT-4o 0.760 63.6    Gemini 1.5 Flash 0.733 79.7 LLMs (different models) 0.576 62.9 2
Table 5: Inter-Annotator Agreement in different annotator groups on the SummEval benchmark. %agreement is the proportion of pairwise annotator comparisons.
Experts
While it’s true that reconciliation naturally leads to increased agreement, the significant improvement in IAA we observed highlights its importance. Though this phase is less common in practice, it is crucial not only for increasing agreement but also for improving the overall quality of annotations and ensuring more reliable outcomes. Interestingly, label changes in this phase were not symmetric, as most changes (69.3%) were in the direction of consistent $\rightarrow$ inconsistent, where one annotator found an inconsistency that the other did not (see all change details in Figure 11). It is important to note that the $\kappa$ obtained by the experts (both before and after reconciliation) was calculated on a more challenging subset, where the original label differed from the LLM prediction, and should be interpreted with this context in mind. This is reflected in the decrease in $\kappa$ observed for all other annotator groups on this subset.
LLMs
GPT-4 and PaLM2, the better-performing LLMs on this task, show high IAA, with $\kappa=0.706$ and $\kappa=0.75$, respectively, which is similar to the experts’ reported $\kappa$. This suggests a comparable level of variance and quality in annotation, providing further empirical evidence for considering LLMs as annotators. This property adds to previous studies showing LLMs’ quality as surrogates for human preferences (Zheng et al., 2023) or evaluations (Chiang and Lee, 2023).
Figure 8: Distribution of crowd-source annotators. Each example was annotated by 3 workers. Plain segments are unanimous annotation, while dotted segments indicate examples where some annotators labeled as inconsistent, and other as consistent. For example, 19.8% of the examples had two inconsistent annotation, and one consistent annotation.
Crowd-Sourcing.
Crowd workers showed near-random agreement, indicating relatively poor-quality annotations. Figure 8 describes the distribution of annotations by MTurk workers. Only 40.8% of the examples were labeled unanimously, whereas the rest included annotations from both classes. In addition, if aggregating by majority vote, we get that 75.8% of the examples are labeled as consistent, which is far from the original distribution of classes. As mentioned before, even experts may miss a small inconsistency nuance, and finding it requires attention. Even from the subset of examples unanimously labeled as consistent, 37.9% have a label of inconsistent in both original and gold labels, which points to a lack of attention and thoroughness.
SummEval.
Table 5 shows the IAA analysis on the SummEval benchmark. We report Krippendorff’s $\alpha$ (Krippendorff, 1970), a generalization of $\kappa$ to account for ordinal labeling. LLMs exhibit high IAA (compared to experts’ IAA) of $\alpha=0.57$ and 62.9% agreement between models, with high consistency across prompts for the same model. Crowd-workers obtain decent results (maybe due to stricter pre-qualification criteria of 10,000 approved HITs), yet they still fall short compared to experts or LLMs.
B.3 Cost and Scalability
In MTurk platform, a total of $400\times 3=1200$ annotations cost 572\$572\\$, including 2 small pilot experiments. All annotations were prepared within a few hours. However, it demanded an additional and significant time for review, after which rejected examples returned to the pool. This annotation-review cycle was conducted for $\sim 5$ iterations. Inference via OpenAI’s API on GPT-4 cost $\sim 4.5\$$ per prompt. Inference via VertexAI’s API on PaLM2 cost $\sim 0.15\$$ per prompt. Both took $\sim 8$ minutes per prompt. Inference on Mistral and Llama3 was via the HuggingFace API, and its cost is estimated by the cost of using a suitable Virtual Machine (VM) on Google Cloud Platform (GCP) for the time of inference (1 minute per model)- $\sim 0.1\$$ per prompt.
Appendix C Annotation
C.1 Crowd-source
@@
Each example was annotated by three annotators, who in addition to the binary label were requested to provide their confidence in their answer, and also write a short explanation for why they chose this label. Pre-qualifications included 50+ approved HITs and 97%+ approval rate, which are at standard scale for the MTurk platform (Kazai et al., 2013; Hauser et al., 2021; Chmielewski and Kucker, 2019). Also, locations were limited to [USA, UK, Australia], which are all English-speaker countries. We disabled the possibility of right-click and Ctrl+c in the platform (as suggested by Veselovsky et al. (2023a)), to prevent (as much as possible) the case where generative-AI (e.g., ChatGPT) will be applied to solve the task instead of humans solving it themselves (as shown by (Veselovsky et al., 2023b)). The maximum time allowed per HIT was 6 minutes, while the actual average execution time was 2:20 minutes for all assignments, and 3 minutes for approved assignments. The guidelines provided to annotators and the annotation platform layout are presented in Figure 9.
@@
Each annotation was manually reviewed and was rejected if the answers were not in line with the instructions, or if it was obvious that the task was not done honestly. Overall, this task suffered from a high rejection rate of 49.2% (1163 rejected, 1200 approved). The main rejection reasons were: lack of meaningful explanation, obvious copy-paste annotations across different examples, explanations contradicting the label annotation, and cases where the explanation was a copy-paste of either the grounding or the statement.
@@
C.2 Experts
Experts annotation was using the platform of Label Studio.  [^1] https://labelstud.io/ Layout design is presented in Figure 10. Examples were presented in random order, and neither the LLM prediction nor the original label were presented during the annotation. In the first stage, each example was annotated independently by both experts. Afterward, the human experts began in a second phase of a reconciliation, where a discussion was made over examples they disagreed over. This reconciliation phase ended up with a much higher agreement and higher-quality labels. Complete agreement was reached in nearly all cases; only a very small number of examples remained unresolved, which may reflect inherent label variation rather than clear annotation errors (Weber-Genzel et al., 2024).
In the reconciliation phase, we observed that most changes (69.3%) were from label 1 to label 0, indicating that contradictions might be hard to find, and not all annotators catch them at first. For the full distribution of label change in the reconciliation phase, see Figure 11.
Figure 11: How experts’ annotations have changed after the reconciliation phase. Most changes occur from 1 (consistent) to 0 (inconsistent).
C.3 LLMs
To annotate a total of $160\times 4=640$ examples from four different datasets, we used four LLMs: GPT-4 (gpt-4-1106-preview) (OpenAI, 2023), PaLM2 (text-bison@002) (Anil et al., 2023), Mistral (7B) [^2] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) and Llama 3 (8B) [^3] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct (Dubey et al., 2024).
Each model was run with four different prompts (see full prompts in Figure 12). We used a variety of terminology, as this task appears to have different framings in different studies. For example, the premise-hypothesis terminology from classic NLI (MacCartney and Manning, 2009), or document-statement used in (Tam et al., 2023). The ensemble reported in the main text refers to ensembling GPT-4 and PaLM2 over four prompts, while the other models are intended for extending our analysis to more models.
@@ C.3 LLMs
Appendix D Data
For our main experiments, we used the TRUE benchmark for factual consistency. Specifically, we focus on four TRUE datasets, one from each task (summarization, dialogue, fact verification, paraphrasing):
MNBM (Maynez et al., 2020): Summarization.
This dataset provides annotations for hallucinations in generated summaries from the XSum dataset (Narayan et al., 2018). Grounding refers to the source document that the summary is based on, while Generated Text consists of model-generated summaries, which may include hallucinated information not present in the source. Three human annotators, trained for the task through two pilot studies, annotated the dataset for the existence of hallucinations. In TRUE, the binary annotations were determined by majority vote.
BEGIN (Dziri et al., 2022): Dialogue.
This dataset evaluates groundedness in knowledge-grounded dialogue systems, where responses are expected to align with an external Grounding source, typically a span from Wikipedia. Generated Text refers to model-generated dialogue responses that were fine-tuned on datasets like Wizard of Wikipedia (Dinan et al., 2019). Data was annotated into entailment/neutral/contradiction labels, by three human annotators, trained for the task through two pilot studies, aggregated by majority vote. In TRUE, binary annotations were then determined by the entailment/not-entailment partition.
VitaminC (Schuster et al., 2021): Fact Verification.
This dataset is based on factual revisions of Wikipedia. The evidence, or Grounding, consists of Wikipedia sentences, either before or after these revisions. Most human involvement came from creating Generated Text rather than the annotation process, with annotators writing claim/evidence pairs derived from Wikipedia revisions, inherently generating labeled data for fact verification. Synthetic examples from the FEVER dataset (Thorne et al., 2018) were also included. Additionally, three annotators reviewed 2,000 examples, presumably to ensure data quality.
PAWS (Zhang et al., 2019): Paraphrasing.
This dataset consists of paraphrase and non-paraphrase pairs. Grounding refers to source sentences drawn from Quora and Wikipedia, while Generated Text was automatically generated through controlled word swapping and back-translation. Five human annotators annotated the dataset with binary labels w.r.t paraphrasing correctness. The dataset includes both high- and low-agreement annotations.
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 13: A model-specific analysis: When LLMs disagree with original labels - who is correct? Overall trend holds: as the LLM’s confidence grows, so does the precision of identifying an error in the original labels. To maintain a balanced number of examples per bin, the bin edges from Figure 2 were slightly adjusted per model due to natural variability and calibration differences. For simplicity in the shared plot across all models, we label the bins as bin1 through bin5, where bin1 starts at 0.5 and bin5 ends at 1.0.
Appendix E Model-Specific Experiments
Our main analysis relies on an ensemble-based approach, which abstracts away from individual model behavior and leverages their collective strength. This design improves alignment with expert annotations, reduces variance, and avoids the need for model selection or prompt-specific tuning. As such, it provides a more stable and generalizable signal than any single model. The ensemble results are presented in Figure 2.
For completeness, we also provide a model-specific analysis of the same phenomenon. Figure 13 reports the percentage of cases where experts agreed with the LLM prediction rather than the original label, broken down by confidence bins and shown separately for GPT-4, PaLM2, LLaMA-3, and Mistral. These curves correspond directly to the red bars in Figure 2, but now reveal each model’s contribution.
Across models, we observe the same overall trend: when models express higher confidence in a label that differs from the original annotation, experts are increasingly likely to agree with them. The magnitude and variance of this effect, however, differ by model. Some models, such as LLaMA-3, display clearer calibration, while others, such as Mistral, show flatter patterns.
A single model is cheaper and can capture the main trend, but its behavior varies by model. Our ensemble reduces these differences, yielding smoother calibration (Figure 2), more consistent agreement with experts, and lower variance—benefits we believe justify the extra compute.
[Uncaptioned image][Uncaptioned image]
Appendix F Mislabeled Data Implications
F.1 Fine-tuning
Hardware.